OpenStudy (anonymous):

!!!! PLEASE HELP!!!!! Taking into account reaction time, the distance d (in feet that a car requires to come to a complete stop while traveling r miles per hour is given by the following function. d(r)=6.97r-90.33 a) express the speed r at which the care is traveling as a function of the distance d required to come to a complete stop. verify your answer by checking that r(d(r))=r, d(r(d))=d r(d)=_________ b) predict the speed that a car was traveling if the distance required to stop was 330 feet. The car was traveling approximately _______ miles per hour. (round to the nearest whole number)

OpenStudy (dumbcow):

\[r = \frac{d+90.33}{6.97}\] d = 330 \[r = \frac{330+90.33}{6.97} = 60.3\]

OpenStudy (anonymous):

Thank you SUPER MUCH!